OpenStudy (anonymous):

Find the equation of a line A(-5,2) B(0,6)

OpenStudy (anonymous):

the slope=(6-2)/(0-(-5))=4/5 so y=4x/5+c for x=0, y=6 6=4*0/5+c 6=0+c c=6 y=4x/5+6 5y=4x+30 5y-4x-30=0